OpenStudy (anonymous):

Find the slope of the line passing through the points (-2, 9) and (5, 3).

OpenStudy (anonymous):

Find the slope of the line passing through the points (-2, 9) and (5, 3). Coordinate 1 = (-2, 9) Coordinate 2 = (5, 3) Coordinates are written as (x,y) m = (y1 - y2) / (x1 - x2) m = (9 - 3) / (-2 - 5) m = 6 / -7

OpenStudy (anonymous):

no answer mine

OpenStudy (anonymous):

now

OpenStudy (anonymous):

or i press delete

OpenStudy (anonymous):

come on

OpenStudy (anonymous):

okay i am hold on

OpenStudy (anonymous):

weres yours??